### v1

@@ 5 Results
In the previous sections, we have presented the new extensions to the physics included in GENEC to allow the evolution of stellar models in the high density/temperature regimes typically found in the last stages of evolution of massive stars. We have thus employed this new version to produce a small sample of models that are run to the pre-supernova stage. The choice of models studied are a 15 solar mass model at solar metallicity, 20 and 25 solar mass models, both run at Large Magellanic Cloud (LMC) metallicity. To compare the results with the current state of the art, we also run the same models with MESA (Paxton et al. 2011, 2013, 2018, 2019) using version 10398. The physics of the GENEC models not explicitly mentioned in this paper are the same as in Ekström et al. (2012). We point out, for example, the mass loss prescription and the mixing as two major physical ingredients that are unchanged. For the MESA models, we use convective boundary mixing (CBM) with a value of $f=0.02$ at the top of the convective layer and $f=0.004$ below. This induces slightly more mixing than the prescription of GENEC. Thus, the MESA models are expected to behave as slightly more massive stars than the GENEC models with the same initial mass. The Schwarzschild criterion defines the convective boundaries in our model, and as such, we did not need to implement semi-convective mixing. The MESA models are run until the pre-SN link, defined as the point where the maximum infall velocity reaches $10^{8}\,\text{cm\,s}^{-1}$. For the GENEC models the acceleration term required to track the in-fall velocity is not implemented yet and, thereby, we do not share a similar definition of the stopping point. To maintain a consistent stopping point, however, we use the central temperature of the MESA model one second before core collapse, and take the first GENEC model with the same core temperature to define this as our pre-supernova model. This ensures that we do not evolve the GENEC models further than they should go due to the lack of core contraction (supplied by the acceleration term), and also that the last GENEC model is extremely close to core collapse. Hence, this model can be followed up by hydrodynamic codes with ease.
@@ 5.1 Core evolution
In this section, we focus on the evolution of central conditions for all models, as most of the novel improvements in GENEC affect the high temperature and high density conditions generally reached in the inner core.
@@
Figure 6: The evolution of the central value of the electron fraction for all models. The cyan model displays the $Y_{e}$ evolution of a $15M_{\odot}$ MESA model with 10 times slower electron capture rates. With horizontal dot-dashed lines, we show the value of $Y_{e}$ that the centre of the star would attain if it were composed of pure $\rm{}^{56}Fe$ (gray) and $\rm{}^{56}Cr$ (magenta). The $\rm{}^{28}Si$ core burning begins, approximately, when $Y_{e}$ decreases from 0.5. We indicate the interval of the end of Si core burning by the magenta region.
@@
In Fig. 5, we compare the evolution of central temperature versus central density for our different models. In the left panel, we separate by the red dotted lines the region where the D-EoS is applied and where the T-EoS is applied, as defined by Eq. (1). In the right panel, we zoom in on the later stages of evolution, where the differences between models are largest. The general aspect of the tracks looks qualitatively (and quantitatively) like the ones shown in Fig. 1 of Heger et al. (2001a), computed with the KEPLER stellar evolution code. Comparing the GENEC models to the MESA ones, here we see that most of the same aspects are present in each track, like the various loops and bends occurring at the start of consecutive burning stages in the core. However, the size of the loops and position of bends are not exactly the same from code to code. These differences generally stem from slightly different treatments of convection in the core and in the layers above it.
@@
In the GENEC models, the convective core size is determined firstly by using the Schwarzschild criterion. Then core overshooting is applied only for the hydrogen and helium burning phases where the size of the convective core is increased by $\alpha H_{p}$, with $H_{p}$ being the pressure scale height and $\alpha=0.1$ as in Ekström et al. (2012). Note that this calibration was made to reproduce the Main Sequence observations of stars in the mass range $1.35M_{\odot}<M<9M_{\odot}$, and could require further fine-tuning for the stars considered here. We choose overshooting and convection criterion in our MESA models to give similar behaviours for the treatment of convection in the initial burning phases. In the advanced burning stages, overshooting and undershooting is applied in the MESA models, which will likely contribute to differences between the codes. However, this choice was made to remain consistent with the MESA state-of-the-art.
@@
Locally, the position in the $(\rho_{\rm c},T_{\rm c})$ plane can differ by $\sim 35\%$ among equivalent GENEC and MESA models. The overall core dynamics, marked by the ascension along a path where $\rho_{\rm c}\propto T_{\rm c}^{3}$, brings all models to relatively similar (though not equal) pre-supernova conditions. During the very final approach to collapse, the last straight path in the profiles before reaching $T\approx 10^{9.9}$ K shows a very similar slope. In the inset of Fig. 5 we show the very final contraction of the stellar core before collapse. The evolution of MESA models and GENEC models share the same slope, but with slightly different central conditions, pointing towards a mostly adiabatic contraction detached from the conditions of the rest of the star. In Fig. 5 we place markers for the MESA models indicating time before collapse of 1000s (circle), 100s (square), 10s (cross) and finally 1s (star). Along this final branch, the changes in the core are extremely rapid. The final model of GENEC is chosen to match the central temperature of the MESA model that is one second away from collapse, i.e. marked by the star in Fig. 5. At the very tip of these evolution tracks, we see a slightly faster contraction for MESA models, which is indicative of the acceleration effects becoming dominant and the collapsing of the core. This is shown clearly in Fig. 5 by the magenta dashed-dotted curve, which is the same $15M_{\odot}$ mass model run in MESA but without using the acceleration term in the stellar evolution equations. The curve of this model and of the other $15M_{\odot}$ MESA model are identical for most of the evolution — we only display the very end phase for this reason — excepted the last second before collapse. Indeed, past the star marker we see the model with the acceleration term contracts slightly faster, i.e., the central density rises faster for a given temperature increase, whereas the model without the acceleration term stays on an adiabatic contraction. This indicates that the absence of the acceleration term in GENEC does not affect the evolution until approximately the last second before collapse, at which point a transition towards fully coupled hydro-codes is preferable to follow the collapse anyway.
@@
The electron fraction evolution of the core leading up to collapse is shown in Fig. 6. The effect of electron captures in our reduced network only becomes relevant once core $\rm{}^{28}Si$ burning begins. The corresponding decrease in the electron fraction of the core is followed in our simplified GeValNet25 network by the electron capture chain, detailed in Sec. 4. A large drop in $Y_{e}$ generally occurs by the end of $\rm{}^{28}Si$ core burning, where the inner core is entirely made up of iron group elements and as such susceptible to electron capture. In Fig. 6 we highlight this region by a magenta band showing roughly the depletion of $\rm{}^{28}Si$ in the core. The electron fraction evolution flattens as the model approaches collapse, as expected due to the quasi beta equilibrium that is reached in the core (Heger et al. 2001b). The final value of $Y_{e}$ is heavily dependent on the mixture of $\rm{}^{56}Fe$ and $\rm{}^{56}Cr$. Therefore, the precise value of the central $Y_{e}$ obtained in these models is closely related to the rates used for the modelled reaction $\rm{}^{56}Fe\ \xrightarrow{(2)}\ \rm^{56}Cr$. This chain is only a rough approximation for how the electron fraction changes due to weak reactions. In this work we intend only to reproduce the basic functionality of the reduced MESA approx21_plus_co56 network. The current network aims only to reproduce the basic properties found in the pre-collapse phase of massive stars due to neutronisation. The final $Y_{e}$ reached by the GENEC models in the core differs from that of MESA, but are all within the range of larger and more sophisticated networks, (see e.g. Farmer et al. 2016; Heger et al. 2001a). The evolution of $Y_{e}$ and the values reached before core collapse are very similar for GENEC and MESA for models with masses $\geq 20\,M_{\odot}$, while there is some moderate discrepancy for the smaller mass under consideration ($15\,M_{\odot}$). To further understand this discrepancy, the cyan dash-dotted line in Fig. 6 shows the same $15M_{\odot}$ MESA model but run with electron captures that are 10 times slower between $\rm{}^{56}Fe$ and $\rm{}^{56}Cr$. We remind here that in reduced-type networks, the speed of the final part of the electron capture chain is calibrated by an artificial factor. In the reduced MESA approx21_plus_co56 network, this is by default $10^{-4}$. The extra model shown in Fig. 6 is run with $10^{-5}$, which is the same artificial factor that the GENEC models are run with. We see that slowing down the final step of the electron capture chain, the $Y_{e}$ evolution becomes much more similar to that of the $15M_{\odot}$ GENEC. This demonstrates some sensitivity of the central value, and thus the core structure, of $Y_{e}$ to this factor. Furthermore, using a single value for this calibration factor over the entire range of massive stars is most likely not adequate to imitate the real behaviour of larger electron capture networks. It is clear that in both reduced nuclear networks (GENEC and MESA) the treatment of electron captures is approximate, and these captures tend to be most important in the low-mass end of the massive star distribution. Therefore, approximating with the same recipe over all masses may lead to discrepancies. A more complete electron capture network (optimally including also beta decays) is required to accurately track the core neutronisation.
@@
Figure 7: Evolution for all models of the Chandrasekhar mass, in solid line, computed using Eq. 18. Also shown is the iron core mass, dotted line, defined by the mass coordinate where the sum of elements heavier than $\rm{}^{48}Cr$ goes below 0.5. Furthermore, we show the degenerate iron core mass that we defined in Eq. 19. In the left panel are shown the GENEC models and in the right the MESA models.
@@
Figure 8: Entropy per baryon (red) and the degeneracy parameter $\Psi$ (black), as a function of mass for the pre-collapse models (see Appendix A, for details on the degeneracy parameter). Solid and dashed-dotted lines correspond to GENEC and MESA models, respectively. From left to right, we show the $15M_{\odot},20M_{\odot}$, and $25M_{\odot}$ models.
@@
In Baron & Cooperstein (1990) the authors investigate the role of this critical mass and add different corrections to the simplified estimation of Eq. 17 written as,
@@
where $S_{e}$ is the entropy of electrons per nucleon and $\rm A$ the mass fraction weighted atomic number. For the iron core, we define it in the same manner as is done in the massive star evolution literature, (see e.g. Hirschi et al. 2004; Chieffi et al. 1998; Paxton et al. 2011; Heger et al. 2001b). The edge of the iron core is taken where the sum of mass fractions of all elements heavier than ${}^{48}\rm Cr$ goes below 0.5, and we denote the enclosed mass by $M_{\rm Fe}$. However, for the purposes of comparing this mass with the Chandrasekhar mass, we also consider the degeneracy of the matter contained within the iron core. Indeed, the original Chandrasekhar computation is done in the context of fully degenerate matter, which is not necessarily the case for what most authors define as the iron core. Therefore, we compute here $M_{\rm Fe,deg}$, defined by
@@
where $f(\Psi)$ is a weighting function varying from 0 to 1 with the following expression,
@@ 5.2 Pre-collapse structure
As we have seen in the previous section, the very final hours of evolution in both GENEC and MESA models will lead to a core collapse due to neutronisation of the core and reaching the limits of instability with respect to the Chandrasekhar mass of the core. In its current version, GENEC does not include the acceleration term in the mechanical equilibrium equation. Without this term, we cannot define a pre-SN model stopping point (the so-called pre-SN link), as is commonly done in the literature, when the maximum in-fall velocity reaches 1000 $\rm km\,s^{-1}$. As explained in the previous sections, we take as pre-SN link definition for the GENEC models the profile where the central temperature is the same as the MESA model equivalent one second before reaching the 1000 $\rm km\,s^{-1}$ infall velocity limit. This is the model shown by the star in the inset of Fig. 5. In this section, we compare the structure of the final models produced using GENEC and MESA and highlight the interesting properties displayed for supernova progenitors. A summary of the characteristics discussed in this section is provided in Table 2.
@@
In Fig. 8 we show the profiles of entropy and the degeneracy parameter $\Psi$ for the final models. The entropy within the inner 1 $M_{\odot}$ is almost identical between codes, large differences appear in the outer regions where the exact boundaries between burning shells, and the strength of mixing, plays an important role. The general trend of degeneracy is coherent between codes, and goes as expected, i.e., the 15 $M_{\odot}$ being the most degenerate and the 25 $M_{\odot}$ the least. The 15 $M_{\odot}$ model of MESA is more degenerate in the core, which can also be seen in Fig. 5 where the MESA model reaches higher densities for a given temperature. We note that the shell structure beyond the inner core of roughly $1\,M_{\odot}$, differs between codes. See for example the two steps in the entropy for the GENEC $25\,M_{\odot}$ at $M_{r}=2M_{\odot}$ and then at $M_{r}=2.5M_{\odot}$, whereas the MESA model only shows one step. Similar differences can be observed in the other models. These stem from different evolution of the various shell burning events. They can be due to variations in the nuclear reaction network, which use different rates and compute different energy generation values, or also differences in the treatment of convection and overshooting. The main differences between the models at collapse, between different codes, is not in the central values of physical quantities, but rather in the stellar structure which can have large consequences for the type of collapse and explosion that may ensue.
@@
In Fig. 9 we show the chemical composition for all pre-supernova models. There are numerous differences both between codes and between models of different masses. We point out the generally smaller C/O core for GENEC models, i.e. the helium shell begins at lower masses in GENEC than in MESA (see also Tab. 2). This is caused by overshooting differences in the earlier phases. The $\rm{}^{28}Si$ shell for the GENEC 20 $M_{\odot}$ model disappears, with a part going to the iron core and the rest merging with the Oxygen shell. This happens in no other models, however, the merging of the Si/O shell has begun in the MESA 15 $M_{\odot}$ where we see silicon being mixed out until $M_{r}\approx 3M_{\odot}$. With the simplistic network used here, the inner core composition is dominated by $\rm{}^{56}Fe$ and $\rm{}^{56}Cr$. The relatives abundances of these two species mostly shapes the $Y_{e}$ profile in the core. Furthermore, we indicate by the light blue hashed regions where convection is active in the star in Fig. 9, the active convection regions vary significantly in all models and provide no clear rule of predicting at collapse what shells are unstable to convection.
@@
Between codes, the structure of the silicon/oxygen shells at the point of collapse can vary quite significantly. As in Yoshida et al. (2019), we define the quantity $f_{\rm Si,O}$,
@@
so that $f_{\rm Si,O}=1$ if the whole region from $r_{8}=1$ to $r_{8}=10$ has $X_{{}^{16}\rm O}>0.1$ and $X_{{}^{28}\rm Si}>0.1$. This helps quantify the amount of potential turbulent mixing in the silicon/oxygen shell at the pre-supernova stage. When such mixing occurs as the collapse of the core begins can lead to impacting the properties of the subsequent explosion. It also may indicate the need for a better follow up of the model in multi-dimension simulations prior to collapse (see Sec. 6). The values for the pre-supernova models are given in Tab. 2. Note that the higher the value of $f_{\rm Si,O}$, the larger the overlapping region between oxygen and silicon shells. In practice, if these shells are overlapping, a merger event is occurring. Hence, a value of $f_{\rm Si,O}$ close to 1 is indicative of the potential for strong turbulent mixing in the Si/O shells. The speed of this mixing can be very large, in Fig. 10 we show the value of the convective Mach number in such a shell reaching similar values as multidimensional studies such as Varma & Müller (2021) and Yoshida et al. (2021) where the turbulent velocity in the Si/O rich layer was reaching several $100\,\rm km\,s^{-1}$ causing very rapid mixing with convective turnover timescales in these shells of order 10’s of seconds.

### v2

@@ 5 Results
In the previous sections, we have presented the new extensions to the physics included in GENEC to allow the evolution of stellar models in the high density and high-temperature regimes typically found in the last stages of evolution of massive stars. We have thus employed this new version to produce a small sample of models that are run to the pre-supernova stage. The choice of models studied are a 15 solar mass model at solar metallicity, 20 and 25 solar mass models, both run at Large Magellanic Cloud (LMC) metallicity. To compare the results with the current state of the art, we also run the same models with MESA (Paxton et al. 2011, 2013, 2018, 2019) using version 10398. The physics of the GENEC models not explicitly mentioned in this paper are the same as in Ekström et al. (2012). We point out, for example, the mass loss prescription and the mixing as two major physical ingredients that are unchanged. For the MESA models, we use convective boundary mixing (CBM) with a value of $f=0.02$ at the top of the convective layer and $f=0.004$ below. This induces slightly more mixing than the prescription of GENEC. Thus, the MESA models are expected to behave as slightly more massive stars than the GENEC models with the same initial mass. The Schwarzschild criterion defines the convective boundaries in our model, and as such, we did not need to implement semi-convective mixing. The MESA models are run until the pre-SN link, defined as the point where the maximum infall velocity reaches $10^{8}\,\text{cm\,s}^{-1}$. For the GENEC models the acceleration term required to track the in-fall velocity is not implemented yet and, thereby, we do not share a similar definition of the stopping point. To maintain a consistent stopping point, however, we use the central temperature of the MESA model one second before core collapse, and take the GENEC model that first reaches the same core temperature and define this as our pre-supernova model. This ensures that we do not evolve the GENEC models further than they should go due to the lack of core contraction (supplied by the acceleration term), and also that the defined pre-SN GENEC model is extremely close to core collapse. Hence, this model can be followed up by hydrodynamic codes with ease.
@@ 5.1 Core evolution
In this section, we focus on the evolution of central conditions for all models, as most of the novel improvements in GENEC affect the high-temperature and high-density conditions generally reached in the inner core.
@@
Figure 7: Evolution of the central value of the electron fraction for all models. The cyan model displays the $Y_{e}$ evolution of a $15M_{\odot}$ MESA model with ten times slower electron capture rates. The horizontal dot-dashed lines show the value of $Y_{e}$ that the centre of the star would attain if it were composed of pure $\rm{}^{56}Fe$ (grey) and $\rm{}^{56}Cr$ (magenta). The $\rm{}^{28}Si$ core burning begins, approximately, when $Y_{e}$ decreases from 0.5. We indicate the interval of the end of Si core burning by the magenta region.
@@
In Fig. 6 we compare the evolution of central temperature versus central density for our different models. In the left panel, we separate by the red dotted lines the region where the D-EoS is applied and where the T-EoS is applied, as defined by Eq. (1). We note that the transition across the boundary between the two segments of the EoS in Fig. 6 is free of numerical artifacts, reflecting the smoothness of the connection between the D-EoS and T-EoS under the prescribed matching conditions, we detail further the matching point in Appendix C. In the right panel, we zoom in on the later stages of evolution, where the differences between models are largest. The general aspect of the tracks looks qualitatively (and quantitatively) like the ones shown in Fig. 1 of Heger et al. (2001a), computed with the KEPLER stellar evolution code. Comparing the GENEC models to the MESA ones, here we see that most of the same aspects are present in each track, like the various loops and bends occurring at the start of consecutive burning stages in the core. However, the size of the loops and position of bends are not exactly the same from code to code. These differences generally stem from slightly different treatments of convection in the core and in the layers above it.
@@
In the GENEC models, the convective core size is determined firstly by using the Schwarzschild criterion. Then core overshooting is applied only for the hydrogen and helium burning phases where the size of the convective core is increased by $\alpha H_{p}$, with $H_{p}$ being the pressure scale height and $\alpha=0.1$ as in Ekström et al. (2012). We note that this calibration was made to reproduce the Main Sequence observations of stars in the mass range $1.35M_{\odot}<M<9M_{\odot}$, and could require further fine-tuning for the stars considered here. We choose the overshooting and convection criterion in our MESA models to give similar behaviours for the treatment of convection in the initial burning phases. In the advanced burning stages, overshooting and undershooting is applied in the MESA models, which will likely contribute to differences between the codes. However, this choice was made to remain consistent with the MESA state-of-the-art.
@@
Locally, the position in the $(\rho_{\rm c},T_{\rm c})$ plane can differ by $\sim 35\%$ among equivalent GENEC and MESA models. The overall core dynamics, marked by the ascension along a path where $\rho_{\rm c}\propto T_{\rm c}^{3}$, brings all models to relatively similar (though not equal) pre-supernova conditions. During the very final approach to collapse, the last straight path in the profiles before reaching $T\approx 10^{9.9}$ K shows a very similar slope. In the inset of Fig. 6 we show the very final contraction of the stellar core before collapse. The evolution of MESA models and GENEC models share the same slope, but with slightly different central conditions, pointing towards a mostly adiabatic contraction detached from the conditions of the rest of the star. In Fig. 6 we place markers for the MESA models indicating time before collapse of 1000s (circle), 100s (square), 10s (cross) and finally 1s (star). Along this final branch, the changes in the core are extremely rapid. The final model of GENEC is chosen to match the central temperature of the MESA model that is one second away from collapse, i.e. marked by the star in Fig. 6. At the very tip of these evolution tracks, we see a slightly faster contraction for MESA models, which is indicative of the acceleration effects becoming dominant and the collapsing of the core. This is shown clearly in Fig. 6 by the magenta dashed-dotted curve, which is the same $15M_{\odot}$ mass model run in MESA but without using the acceleration term in the stellar evolution equations. The curve of this model and of the other $15M_{\odot}$ MESA model are identical for most of the evolution — we only display the very end phase for this reason — excepted the last second before collapse. Past the star marker we see the model with the acceleration term contracts slightly faster (i.e. the central density rises faster for a given temperature increase), whereas the model without the acceleration term stays on an adiabatic contraction. This indicates that the absence of the acceleration term in GENEC does not affect the evolution until approximately the last second before collapse, at which point a transition towards fully coupled hydro-codes is preferable to follow the collapse anyway.
@@
The electron fraction evolution of the core leading up to collapse is shown in Fig. 7. The effect of electron capture in our reduced network only becomes relevant once core $\rm{}^{28}Si$ burning begins. The corresponding decrease in the electron fraction of the core is followed in our simplified GeValNet25 network by the electron capture chain, detailed in Sect. 4. A large drop in $Y_{e}$ generally occurs by the end of $\rm{}^{28}Si$ core burning, where the inner core is entirely made up of iron group elements and as such susceptible to electron capture. In Fig. 7 we highlight this region by a magenta band showing roughly the depletion of $\rm{}^{28}Si$ in the core. The electron fraction evolution flattens as the model approaches collapse, as expected due to the quasi beta equilibrium that is reached in the core (Heger et al. 2001b). The final value of $Y_{e}$ is heavily dependent on the mixture of $\rm{}^{56}Fe$ and $\rm{}^{56}Cr$. Therefore, the precise value of the central $Y_{e}$ obtained in these models is closely related to the rates used for the modelled reaction $\rm{}^{56}Fe\ \xrightarrow{(2)}\ \rm^{56}Cr$. This chain is only a rough approximation for how the electron fraction changes due to weak reactions. In this work we intend only to reproduce the basic functionality of the reduced MESA approx21_plus_co56 network. The current network aims only to reproduce the basic properties found in the pre-collapse phase of massive stars due to neutronisation. The final $Y_{e}$ reached by the GENEC models in the core differs from that of MESA, but are all within the range of larger and more sophisticated networks, (see e.g. Farmer et al. 2016; Heger et al. 2001a). The evolution of $Y_{e}$ and the values reached before core collapse are very similar for GENEC and MESA for models with masses $\geq 20\,M_{\odot}$, while there is some moderate discrepancy for the smaller mass under consideration ($15\,M_{\odot}$). To further understand this discrepancy, the cyan dash-dotted line in Fig. 7 shows the same $15M_{\odot}$ MESA model but run with electron captures that is ten times slower between $\rm{}^{56}Fe$ and $\rm{}^{56}Cr$. We recall here that in reduced-type networks, the speed of the final part of the electron capture chain is calibrated by an artificial factor. In the reduced MESA approx21_plus_co56 network, this is by default $10^{-4}$. The extra model shown in Fig. 7 is run with $10^{-5}$, which is the same artificial factor that the GENEC models are run with. We see that slowing down the final step of the electron capture chain, the $Y_{e}$ evolution becomes much more similar to that of the $15M_{\odot}$ GENEC. This demonstrates some sensitivity of the central value, and thus the core structure, of $Y_{e}$ to this factor. Furthermore, using a single value for this calibration factor over the entire range of massive stars is most likely not adequate to imitate the real behaviour of larger electron capture networks. It is clear that in both reduced nuclear networks (GENEC and MESA) the treatment of electron captures is approximate, and these captures tend to be most important in the low-mass end of the massive star distribution. Therefore, approximating with the same recipe over all masses may lead to discrepancies. A more complete electron capture network (optimally including also beta decays) is required to accurately track the core neutronisation.
@@
Figure 8: Evolution for all models of the Chandrasekhar mass ( solid lines), computed using Eq. 19. Also shown is the iron core mass (dotted line) defined by the mass coordinate where the sum of elements heavier than $\rm{}^{48}Cr$ goes below 0.5. Furthermore, we show the degenerate iron core mass that we defined in Eq. 20. In the left panel are shown the GENEC models and in the right the MESA models.
@@
Figure 9: Entropy per baryon (red) and the degeneracy parameter $\Psi$ (black), as a function of mass for the pre-collapse models (see Appendix B for details on the degeneracy parameter). The solid and dash-dotted lines correspond to the GENEC and MESA models, respectively. From left to right, we show the $15M_{\odot},20M_{\odot}$, and $25M_{\odot}$ models.
@@
In Baron & Cooperstein (1990) the authors investigate the role of this critical mass and add different corrections to the simplified estimation of Eq. 18 written as
@@
where $S_{e}$ is the entropy of electrons per nucleon and $\rm A$ the mass fraction weighted atomic number. For the iron core, we define it in the same manner as is done in the massive star evolution literature, (see e.g. Hirschi et al. 2004; Chieffi et al. 1998; Paxton et al. 2011; Heger et al. 2001b). The edge of the iron core is taken where the sum of mass fractions of all elements heavier than ${}^{48}\rm Cr$ goes below 0.5, and we denote the enclosed mass by $M_{\rm Fe}$. However, for the purposes of comparing this mass with the Chandrasekhar mass, we also consider the degeneracy of the matter contained within the iron core. The original Chandrasekhar computation is done in the context of fully degenerate matter, which is not necessarily the case for what most authors define as the iron core. Therefore, we compute here $M_{\rm Fe,deg}$, defined by
@@
where $f(\Psi)$ is a weighting function varying from 0 to 1 with the expression,
@@ 5.2 Pre-collapse structure
As we have seen in the previous section, the very final hours of evolution in both GENEC and MESA models will lead to a core collapse due to neutronisation of the core and reaching the limits of instability with respect to the Chandrasekhar mass of the core. In its current version, GENEC does not include the acceleration term in the mechanical equilibrium equation. Without this term, we cannot define a pre-SN model stopping point (the pre-SN link), as is commonly done in the literature, when the maximum in-fall velocity reaches 1000 $\rm km\,s^{-1}$. As explained in the previous sections, we take as pre-SN link definition for the GENEC models the profile where the central temperature is the same as the MESA model equivalent one second before reaching the 1000 $\rm km\,s^{-1}$ infall velocity limit. This is the model shown by the star in the inset of Fig. 6. In this section, we compare the structure of the final models produced using GENEC and MESA and highlight the interesting properties displayed for supernova progenitors. A summary of the characteristics discussed in this section is provided in Table 9.
@@
In Fig. 9 we show the profiles of entropy and the degeneracy parameter $\Psi$ for the final models. The entropy within the inner 1 $M_{\odot}$ is almost identical between codes, large differences appear in the outer regions where the exact boundaries between the burning shells, and the strength of mixing, plays an important role. The general trend of degeneracy is consistent between codes, and goes as expected, i.e. the 15 $M_{\odot}$ being the most degenerate and the 25 $M_{\odot}$ the least. The 15 $M_{\odot}$ model of MESA is more degenerate in the core, which can also be seen in Fig. 6 where the MESA model reaches higher densities for a given temperature. We note that the shell structure beyond the inner core of roughly $1\,M_{\odot}$, differs between codes. For example, there are two steps in the entropy for the GENEC $25\,M_{\odot}$ at $M_{r}=2M_{\odot}$ and then at $M_{r}=2.5M_{\odot}$, whereas the MESA model only shows one step. Similar differences can be observed in the other models. These stem from different evolution of the various shell burning events. They can be due to variations in the nuclear reaction network, which use different rates and compute different energy generation values, or also differences in the treatment of convection and overshooting. The main differences between the models at collapse, between different codes, is not in the central values of physical quantities, but rather in the stellar structure which can have large consequences for the type of collapse and explosion that may ensue.
@@
In Fig. 10 we show the chemical composition for all pre-supernova models. There are numerous differences both between codes and between models of different masses. We point out the generally smaller C/O core for GENEC models (i.e. the helium shell begins at lower masses in GENEC than in MESA; see also Table. 9). This is caused by overshooting differences in the earlier phases. The $\rm{}^{28}Si$ shell for the GENEC 20 $M_{\odot}$ model disappears, with a part going to the iron core and the rest merging with the oxygen shell. This happens in no other models; however, the merging of the Si/O shell has begun in the MESA 15 $M_{\odot}$ where we see silicon being mixed out until $M_{r}\approx 3M_{\odot}$. With the simplistic network used here, the inner core composition is dominated by $\rm{}^{56}Fe$ and $\rm{}^{56}Cr$. The relatives abundances of these two species mostly shapes the $Y_{e}$ profile in the core. We may point out here that the MESA models show a higher helium abundance in the core than the GENEC models. This is due to the GeValNet25 implementation not including the necessary reactions to link the QSE group of neutrons, protons, and helium needed in the very final phase. This omission speeds up significantly the nuclear calculations and does not alter significantly the evolution of $Y_{e}$ (which is, anyway, only approximate in our code because of the effective treatment of the weak reactions). Furthermore, we indicate by the light blue hashed regions where convection is active in the star in Fig. 10, the active convection regions vary significantly in all models and provide no clear rule of predicting at collapse what shells are unstable to convection.
@@
Between codes, the structure of the silicon and oxygen shells at the point of collapse can vary quite significantly. As in Yoshida et al. (2019), we define the quantity $f_{\rm Si,O}$,
@@
so that $f_{\rm Si,O}=1$ if the whole region from $r_{8}=1$ to $r_{8}=10$ has $X_{{}^{16}\rm O}>0.1$ and $X_{{}^{28}\rm Si}>0.1$. This helps quantify the amount of potential turbulent mixing in the silicon-oxygen shell at the pre-supernova stage. When such mixing occurs as the collapse of the core begins can lead to impacting the properties of the subsequent explosion. It also may indicate the need for a better follow up of the model in multi-dimension simulations prior to collapse (see Sect. 6). The values for the pre-supernova models are given in Table. 9. We note that the higher the value of $f_{\rm Si,O}$, the larger the overlapping region between oxygen and silicon shells. In practice, if these shells are overlapping, a merger event is occurring. Hence, a value of $f_{\rm Si,O}$ close to 1 is indicative of the potential for strong turbulent mixing in the Si/O shell. The speed of this mixing can be very large, in Fig. 11 we show the value of the convective Mach number in such a shell reaching similar values as multidimensional studies such as Varma & Müller (2021) and Yoshida et al. (2021) where the turbulent velocity in the Si/O rich layer was reaching several $100\,\rm km\,s^{-1}$ causing very rapid mixing with convective turnover timescales in these shells of order 10’s of seconds.